OpenStudy (anonymous):

Show that for any \(n \ge 1\) \[ \frac{1}{2}\le\sum _{r=1}^n \frac{r}{n^2+r}\le\frac{n+1}{2 n} \]

OpenStudy (anonymous):

\[\sum_{r=1}^{n} \frac{r}{n^2+r} \le \sum_{r=1}^{n} \frac{r}{n^2}=\frac{1}{n^2}\sum_{r=1}^{n}r=\frac{\frac{n(n+1)}{2}}{n^2}=\frac{n+1}{2n}\]

OpenStudy (anonymous):

Great. How about the other inequality?

OpenStudy (anonymous):

\[\sum_{r=1}^{n} \frac{r}{n^2+r}\ge \sum_{r=1}^{n} \frac{r}{n^2+n}=\frac{1}{2}\]